# Group Order is in $\mathrm{QCMA}$

### v1

@@ 1.3 Overview of the proof strategy
In our protocol for Group Order Verification (described in Section 5), an honest Merlin sends a decomposition satisfying ()($\star$)( ⋆ ). In order to check if $H_{i}/H_{i-1}\cong S_{i}$ for some specific simple group $S_{i}$, it is thus enough to check if $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}\cong S_{i}$. As explained above, to use the homomorphism test, we need to be able to efficiently check membership in $H_{0}$. Crucially, $H_{0}$ is now a solvable group, and we can thus use Watrous’ polynomial-time quantum algorithm for membership testing in solvable groups [42] to implement the homomorphism test efficiently.
@@ 2 Preliminaries
In this section, we describe the notions of complexity theory, group theory and black-box groups needed to show our results. For any positive integer $s$, we write $[s]=\{1,\ldots,s\}$.
@@ 2.2 Group theory
(see, e.g., [38, Proposition 5.42]) For any finite group $G$ and any prime power $p^{t}$, $p^{t}$ divides $|G|$ if and only if $G$ has a subgroup of order $p^{t}$.
@@
Any finite group has a composition series (with $s=O(\log|G|)$). While a group may have more than one composition series, the Jordan-Hölder theorem (e.g., [16, Theorem 22]) shows that they have the same length and the same composition factors, up to permutation and isomorphism.
@@
In particular, any composition factor of a solvable group has a polylogarithmic-length presentation. For any group isomorphic to a subgroup of a permutation group of small degree, any composition factor also has a polylogarithmic-length presentation. This result is well-established in the literature (see, e.g., [34, Section 6]), but not explicitly stated. We provide a proof here for completeness.
@@
Suppose $H_{i}/H_{i-1}\cong\mathrm{R}(q)$. It is known that $(q^{3}+1)$ is the least positive integer such that there is an injective homomorphism $\phi:\mathrm{R}(q)\to\mathrm{Sym}(q^{3}+1)$ (see, e.g., [20, Table 4]). Therefore, $|H_{i}/H_{i-1}|=|\mathrm{R}(q)|=q^{3}(q^{3}+1)(q-1)\leq(q^{3}+1)^{3}\leq k^{3}$. Taking the generating set as the set of all elements of $H_{i}/H_{i-1}$ and the set of relations as all the multiplication relations in the multiplication table, we get a $\mathrm{poly}(k)$-length presentation for $H_{i}/H_{i-1}\cong\mathrm{R}(q)$. ∎
@@ 2.3 Black-box groups
  • for each $i\in[s]$, a certificate of the inclusion $g_{i}^{p_{i}}\in\langle g_{1},\ldots,g_{i-1}\rangle$
@@
The prover sends elements $g^{\prime}_{1},\ldots,g^{\prime}_{s}$ of $G$, along with certificates of the equality $\langle g^{\prime}_{1},\ldots,g^{\prime}_{s}\rangle=\langle g_{1},\ldots,g_{t}\rangle$. The verifier checks if the certificates are correct (which guarantees that $\langle g^{\prime}_{1},\ldots,g^{\prime}_{s}\rangle=G$) and checks if each relation in $\mathcal{R}$ holds in $G$ when replacing $\alpha_{i}$ by $g^{\prime}_{i}$ for all $i\in[s]$ (which guarantees [^6] This is a folklore fact (see for instance [16, Section 6.3]). Here is a proof: the map $\varphi(\alpha_{i})=g^{\prime}_{i}$ can be extended into a surjective homomorphism $\varphi\colon S\to\langle g^{\prime}_{1},\ldots,g_{s}^{\prime}\rangle$. We thus have $\langle g^{\prime}_{1},\ldots,g_{s}^{\prime}\rangle\cong S/\ker{(\varphi)}$. that the group $\langle g^{\prime}_{1},\ldots,g_{s}^{\prime}\rangle$ is isomorphic to $S/N$ for some normal subgroup $N$ of $S$).
@@ 3.1 The Babai-Beals filtration
A set of generators of $\mathrm{Pker}(G)$ can be computed in Monte Carlo polynomial time [8, Theorem 1.1]. Moreover, in Monte Carlo polynomial time, we can set up a data structure which allows, for any $g\in G$, to compute $\phi(g)$ in deterministic polynomial time ([8, Corollary 5.2]). Using this data structure, membership in $\mathrm{Pker}(G)$ can checked in deterministic polynomial time (see the discussion after Theorem 1.1 in [8]). The following tasks can also be solved in deterministic polynomial time: [^7] As explained after Theorem 1.1 in [8], this can be done by using the extensive library of polynomial-time algorithms for permutation groups [11, 15, 34, 35]. More specifically, we can use the algorithms from [34, Section 6], [35, Section 3] or [15, Section 3.6].
@@ 3.2 Nice decompositions and their properties
Clearly, $\psi$ is a surjective homomorphism and $H_{i-1}\leq\ker(\psi)$. Since $\ker(\psi)\neq H_{i}$, and $H_{i}/H_{i-1}\cong T_{i}$ is a simple group, $H_{i-1}=\ker(\psi)$. By the first homomorphism theorem of groups, we get $H_{i}/H_{i-1}\cong\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$, i.e., Condition ()($\star$)( ⋆ ) holds. Combined with the fact that each $H_{i}/H_{i-1}$ is simple, this implies Condition (C3) of Definition 3.1. We conclude that $\mathrm{Pker}(G)$ has a nice decomposition satisfying Condition ($\star$). ∎
@@ 4 Testing Isomorphism to a Ree Group of Rank One
ReeIso$(\mathcal{G})$     // $\mathcal{G}$ is a black-box group
Input: $\ast$ a set of generators of a solvable subgroup $L\leq\mathcal{G}$ast two elements $\beta,\gamma\in\mathcal{G}$ such that $L$ is normal in $\langle\beta,\gamma,L\rangle$ast an integer $q$ of the form $3^{2a+1}$ for some $a>0$ Output: yes if $\langle\beta,\gamma,L\rangle/L$ is isomorphic to $\mathrm{R}(q)$; no otherwise
@@ 4.1 Constructive membership for the Ree group in its standard representation
Let $M=2^{\mathrm{poly}(\log q)}$ denote the total number of seeds $\lambda$. The number of pairs $(\lambda,M)$ such that $M\in\mathsf{Valid}(\lambda)$ is $(1-10^{-7})M|\mathrm{R}(q)|$. Note that for each $\lambda\notin\mathsf{Good}$ there are at most $(1-10^{-5})|\mathrm{R}(q)|$ pairs $(\lambda,M)$ with $M\in\mathsf{Valid}(\lambda)$, while for each $\lambda\in\mathsf{Good}$ there are (obviously) at most $|\mathrm{R}(q)|$ pairs $(\lambda,M)$ with $M\in\mathsf{Valid}(\lambda)$. We thus have
$|\mathsf{Good}||\mathrm{R}(q)|+(M-|\mathsf{Good}|)(1-10^{-5})|\mathrm{R}(q)|\geq(1-10^{-7})M|\mathrm{R}(q)|,$
@@
$\frac{|\mathsf{Good}|}{M}\geq\frac{10^{-5}-10^{-7}}{10{{}^{-5}}}=0.99,$
@@ 4      $h_{i}=g_{\lambda}(xr)g_{\lambda}(r^{-1})$ ;
In this case, $\mathsf{IsoCheck}$ outputs “no” at Section 4.2 at least once during the 12 iterations with probability at least $0.71-(9/10)^{12}>$ 0.7.
@@
We actually have $\mathsf{Valid}(\lambda)\neq\mathrm{R}(q)$. The probability that the arguments of the $12\cdot 3+50\cdot 2=136$ calls to the function $g_{\lambda}$ performed by $\mathsf{IsoCheck}$ are all in the set $\mathsf{Valid}(\lambda)$ is nevertheless at least $0.981-132\cdot 10^{-5}>$ 0.98. The overall probability that $\mathsf{IsoCheck}$ outputs “no” is thus at least $0.7-0.01-0.02>2/3$. ∎
@@ 5.2 Checking that the order divides $m$
    for each $i\in[s]$ such that $z_{i}$ corresponds to the standard name of some Ree group of rank one $\mathrm{R}(q)$, the certificate from Theorem 4.1 certifying that $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$ is isomorphic to $\mathrm{R}(q)$
@@
    the product $|H_{0}|\cdot|\mathrm{gr}(z_{1})|\cdots|\mathrm{gr}(z_{s})|\cdot m_{1}\cdot m_{2}$ divides $m$
Item (1) can be implemented using Theorem 2.4 and Corollary 2.1. Item (2) can be implemented in deterministic polynomial time using the efficient procedure for membership in $\mathrm{Pker}(G)$ of Section 3.1. Item (3) is trivial to check. Item (4$\sim$(8) can be checked in deterministic polynomial time from the discussion in Section 2.3 (for (7) and (8) we use Proposition 2.2 and Proposition 2.1, respectively). Item (9) can be checked (with high probability) in quantum polynomial time, from Theorem 4.1. To check Item (10), we just need to compute $|H_{0}|$, which can be done with high probability using Theorem 2.5.
@@
With the above choices, checking Items (1) and (9) succeeds with high probability, while checking Items (2$\sim$(8) always succeed. From Condition ($\star$), we know that $H_{i}/H_{i-1}$ is isomorphic to $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$ for all $i\in[s]$. We thus have
@@
Assume that $|G|$ does not divide $m$. If Item (1) or (9) is not true, Arthur rejects with high probability. If Items (2$\sim$(7) are not all true, Arthur always rejects. In the following, we thus assume that Items (1$\sim$(9) are all true.
@@ 6 Proofs of the Other Results
To show that Proper Subgroup is in $\mathrm{QCMA}$, we follow the strategy of [41, Section 5]. Merlin guesses an element $a\in\langle g_{1},\,\ldots,\,g_{k}\rangle$ such that $a\not\in\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ and membership certificates certifying that $a\in\langle g_{1},\,\ldots,\,g_{k}\rangle$ and $h_{i}\in\langle g_{1},\,\ldots,\,g_{k}\rangle$ for each $i\in[\ell]$. Arthur checks that the membership certificates are correct and checks that $a\not\in\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ using the $\mathrm{QCMA}$ protocol for Group Non-Membership.
@@
Watrous [41, Section 5] showed that Simple Group is in $\mathrm{coQMA}$ by using a quantum proof for Group Non-Membership. From Corollary 1.2, we can conclude that Simple Group is in $\mathrm{coQCMA}$. Theorem 2.1, Proposition 2.1 and Theorem 4.1 together imply that Simple Group is in $\mathrm{QCMA}$. ∎
@@
Watrous [41] showed that Intersection, Centralizer and Maximal Normal Subgroup are in $\mathrm{coQMA}$ by using a quantum proof for Group Non-Membership along with classical proofs for various other properties (see, [41, Section 5]). By Corollary 1.2, Group Non-Membership is in $\mathrm{QCMA}$. This implies that Intersection, Centralizer and Maximal Normal Subgroup are in $\mathrm{coQCMA}$. ∎

### v2

@@ 1.3 Overview of the proof strategy
In our protocol for Group Order Verification (described in Section 5), an honest Merlin sends a decomposition satisfying $(\star)$. In order to check if $H_{i}/H_{i-1}\cong S_{i}$ for some specific simple group $S_{i}$, it is thus enough to check if $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}\cong S_{i}$. As explained above, to use the homomorphism test, we need to be able to efficiently check membership in $H_{0}$. Crucially, $H_{0}$ is now a solvable group, and we can thus use Watrous’ polynomial-time quantum algorithm for membership testing in solvable groups [43] to implement the homomorphism test efficiently.
@@ 2 Preliminaries
In this section, we describe the notions of complexity theory, group theory and black-box groups needed to show our results. For any positive integer $s$, we write $s]=\{1,\ldots,s\}$.
@@ 2.2 Group theory
(see, e.g., [39, Proposition 5.42]) For any finite group $G$ and any prime power $p^{t}$, $p^{t}$ divides $|G|$ if and only if $G$ has a subgroup of order $p^{t}$.
@@
Any finite group has a composition series (with $s=O(\log|G|)$). While a group may have more than one composition series, the Jordan-Hölder theorem (e.g., [17, Theorem 22]) shows that they have the same length and the same composition factors, up to permutation and isomorphism.
@@
In particular, any composition factor of a solvable group has a polylogarithmic-length presentation. For any group isomorphic to a subgroup of a permutation group of small degree, any composition factor also has a polylogarithmic-length presentation. This result is well-established in the literature (see, e.g., [35, Section 6]), but not explicitly stated. We provide a proof here for completeness.
@@
Suppose $H_{i}/H_{i-1}\cong\mathrm{R}(q)$. It is known that $(q^{3}+1)$ is the least positive integer such that there is an injective homomorphism $\phi:\mathrm{R}(q)\to\mathrm{Sym}(q^{3}+1)$ (see, e.g., [21, Table 4]). Therefore, $|H_{i}/H_{i-1}|=|\mathrm{R}(q)|=q^{3}(q^{3}+1)(q-1)\leq(q^{3}+1)^{3}\leq k^{3}$. Taking the generating set as the set of all elements of $H_{i}/H_{i-1}$ and the set of relations as all the multiplication relations in the multiplication table, we get a $\mathrm{poly}(k)$-length presentation for $H_{i}/H_{i-1}\cong\mathrm{R}(q)$. ∎
@@ 2.3 Black-box groups
  • for each $i\in[s]$, a certificate of the inclusion $g_{i}^{p_{i}}\in\langle g_{1},\ldots,g_{i-1}\rangle$ .
@@
The prover sends elements $g^{\prime}_{1},\ldots,g^{\prime}_{s}$ of $G$, along with certificates of the equality $\langle g^{\prime}_{1},\ldots,g^{\prime}_{s}\rangle=\langle g_{1},\ldots,g_{t}\rangle$. The verifier checks if the certificates are correct (which guarantees that $\langle g^{\prime}_{1},\ldots,g^{\prime}_{s}\rangle=G$) and checks if each relation in $\mathcal{R}$ holds in $G$ when replacing $\alpha_{i}$ by $g^{\prime}_{i}$ for all $i\in[s]$ (which guarantees [^6] This is a folklore fact (see for instance [17, Section 6.3]). Here is a proof: the map $\varphi(\alpha_{i})=g^{\prime}_{i}$ can be extended into a surjective homomorphism $\varphi\colon S\to\langle g^{\prime}_{1},\ldots,g_{s}^{\prime}\rangle$. We thus have $\langle g^{\prime}_{1},\ldots,g_{s}^{\prime}\rangle\cong S/\ker{(\varphi)}$. that the group $\langle g^{\prime}_{1},\ldots,g_{s}^{\prime}\rangle$ is isomorphic to $S/N$ for some normal subgroup $N$ of $S$).
@@ 3.1 The Babai-Beals filtration
A set of generators of $\mathrm{Pker}(G)$ can be computed in Monte Carlo polynomial time [8, Theorem 1.1]. [^7] Theorem 1.1 in [8] assumes that a superset of the set of primes dividing $|G|$ is given as an additional input. These primes are used to compute the order of elements of $G$ efficiently. In the quantum setting, this is not needed since the order of an element can be computed in polynomial time [15, 40]. Moreover, in Monte Carlo polynomial time, we can set up a data structure which allows, for any $g\in G$, to compute $\phi(g)$ in deterministic polynomial time ([8, Corollary 5.2]). Using this data structure, membership in $\mathrm{Pker}(G)$ can checked in deterministic polynomial time (see the discussion after Theorem 1.1 in [8]). The following tasks can also be solved in deterministic polynomial time: [^8] As explained after Theorem 1.1 in [8], this can be done by using the extensive library of polynomial-time algorithms for permutation groups [11, 16, 35, 36]. More specifically, we can use the algorithms from [35, Section 6], [36, Section 3] or [16, Section 3.6].
@@ 3.2 Nice decompositions and their properties
Clearly, $\psi$ is a surjective homomorphism and $H_{i-1}\leq\ker(\psi)$. Since $\ker(\psi)\neq H_{i}$, and $H_{i}/H_{i-1}\cong T_{i}$ is a simple group, $H_{i-1}=\ker(\psi)$. By the first homomorphism theorem of groups, we get $H_{i}/H_{i-1}\cong\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$, i.e., Condition $(\star)$ holds. Combined with the fact that each $H_{i}/H_{i-1}$ is simple, this implies Condition (C3) of Definition 3.1. We conclude that $\mathrm{Pker}(G)$ has a nice decomposition satisfying Condition ($\star$). ∎
@@ 4 Testing Isomorphism to a Ree Group of Rank One
[…]
Input: $\ast$ a set of generators of a solvable subgroup $L\leq\mathcal{G}$ $\ast$ two elements $\beta,\gamma\in\mathcal{G}$ such that $L$ is normal in $\langle\beta,\gamma,L\rangle$ $\ast$ an integer $q$ of the form $3^{2a+1}$ for some $a>0$ Output: yes if $\langle\beta,\gamma,L\rangle/L$ is isomorphic to $\mathrm{R}(q)$; no otherwise
@@ 4.1 Constructive membership for the Ree group in its standard representation
Let $\delta=2^{\mathrm{poly}(\log q)}$ denote the total number of seeds $\lambda$. The number of pairs $(\lambda,M)$ such that $M\in\mathsf{Valid}(\lambda)$ is $(1-10^{-7})\delta|\mathrm{R}(q)|$. Note that for each $\lambda\notin\mathsf{Good}$ there are at most $(1-10^{-5})|\mathrm{R}(q)|$ pairs $(\lambda,M)$ with $M\in\mathsf{Valid}(\lambda)$, while for each $\lambda\in\mathsf{Good}$ there are (obviously) at most $|\mathrm{R}(q)|$ pairs $(\lambda,M)$ with $M\in\mathsf{Valid}(\lambda)$. We thus have
$|\mathsf{Good}||\mathrm{R}(q)|+(\delta-|\mathsf{Good}|)(1-10^{-5})|\mathrm{R}(q)|\geq(1-10^{-7})\delta|\mathrm{R}(q)|,$
@@
$\frac{|\mathsf{Good}|}{\delta}\geq\frac{10^{-5}-10^{-7}}{10{{}^{-5}}}=0.99,$
@@ 4 $h_{i}=g_{\lambda}(xr)g_{\lambda}(r^{-1})$ ;
In this case, $\mathsf{IsoCheck}$ outputs “no” at Section 4.2 at least once during the 12 iterations with probability at least $1-(9/10)^{12}>0.7$.
@@
We actually have $\mathsf{Valid}(\lambda)\neq\mathrm{R}(q)$. The probability that the arguments of the $12\cdot 3+50\cdot 2=136$ calls to the function $g_{\lambda}$ performed by $\mathsf{IsoCheck}$ are all in the set $\mathsf{Valid}(\lambda)$ is nevertheless at least $1-132\cdot 10^{-5}>0.98$. The overall probability that $\mathsf{IsoCheck}$ outputs “no” is thus at least 0.70.010.02>2/30.7-0.01-0.02>2/3. ∎
@@
    for each $i\in[s]$ such that $z_{i}$ corresponds to the standard name of some Ree group of rank one $\mathrm{R}(q)$, the certificate from Theorem 4.1 certifying that $\langle H_{0},\beta_{i},\gamma_{i}\rangle/H_{0}$ is isomorphic to $\mathrm{R}(q)$ .
@@
    the product $|H_{0}|\cdot|\mathrm{gr}(z_{1})|\cdots|\mathrm{gr}(z_{s})|\cdot m_{1}\cdot m_{2}$ divides $m$ .
[…]
@@
[…]
@@
[…]
@@ 6 Proofs of the Other Results
To show that Proper Subgroup is in $\mathrm{QCMA}$, we follow the strategy of [42, Section 5]. Merlin guesses an element $a\in\langle g_{1},\,\ldots,\,g_{k}\rangle$ such that $a\not\in\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ and membership certificates certifying that $a\in\langle g_{1},\,\ldots,\,g_{k}\rangle$ and $h_{i}\in\langle g_{1},\,\ldots,\,g_{k}\rangle$ for each $i\in[\ell]$. Arthur checks that the membership certificates are correct and checks that $a\not\in\langle h_{1},\,\ldots,\,h_{\ell}\rangle$ using the $\mathrm{QCMA}$ protocol for Group Non-Membership.
@@
Watrous [42, Section 5] showed that Simple Group is in $\mathrm{coQMA}$ by using a quantum proof for Group Non-Membership. From Corollary 1.2, we can conclude that Simple Group is in $\mathrm{coQCMA}$. Theorem 2.1, Proposition 2.1 and Theorem 4.1 together imply that Simple Group is in $\mathrm{QCMA}$. ∎
@@
Watrous [42] showed that Intersection, Centralizer and Maximal Normal Subgroup are in $\mathrm{coQMA}$ by using a quantum proof for Group Non-Membership along with classical proofs for various other properties (see, [42, Section 5]). By Corollary 1.2, Group Non-Membership is in $\mathrm{QCMA}$. This implies that Intersection, Centralizer and Maximal Normal Subgroup are in $\mathrm{coQCMA}$. ∎